OpenStudy (anonymous):

What are the odds of rolling a number greater than 3 on a six sided die? i got 3:6 or 1:2

OpenStudy (campbell_st):

well you can have 4, 5 or 6 so its 3 chances out of 6 possible

OpenStudy (kropot72):

The odds of rolling 4 are 1/6, rolling 5 are 1/6 and 6 are 1/6. the three outcomes are mutually exclusive. therefore the probabilities are added. \[P(4\ or\ 5\ or\ 6) = \frac{1}{6}+\frac{1}{6}+\frac{1}{6}=\frac{3}{6}=\frac{1}{2}\]

OpenStudy (anonymous):

so yes? :)

OpenStudy (anonymous):

i am correct?

OpenStudy (campbell_st):

yes

Directrix (directrix):

Probability and Odds are not the same although they are closed related. Odds is the ratio of successful outcomes to unsuccessful outcomes. In this case, there are 3 successful outcomes {4, 5, 6} There are 3 unsuccessful outcomes {1, 2, 3}. The odds in favor of rolling a number greater than 3 is 3 to 3 or 1:1. Note: 1/2 or 1:2 is INCORRECT @anthony~